OpenStudy (anonymous):

Which equation is not equivalent to the formula r = st? s equals r over t t equals r over s s equals t over r r = ts

OpenStudy (anonymous):

s=r/t not t/r so the third one is wrong

OpenStudy (anonymous):

you can automatically rule out D now set it equal to each variable:\[r=st\]\[s={r\over t}\]\[t={r\over s}\]

OpenStudy (anonymous):

now out of your choices -A, B, and C- which is not equivalent?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Is it a

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

nope, try again

OpenStudy (anonymous):

s=r/t r=ts t=r/s r=ts s=t/r rs=t

OpenStudy (anonymous):

then its b

OpenStudy (anonymous):

i gave you all the equivalents, tell me which one is not an equivalent

OpenStudy (anonymous):

B says \(t=\large {r\over s}\)..i told you \(t=\large{r\over s}\)

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

r = st r/t = s r/s = t 1. correct 2. correct 3. incorrect 4. correct